OpenStudy (anonymous):

Given the points (2,3) and (6,8), write the point-slope form of the equation. A) y - 3 = x - 2 B) - 3 = 5/4 (x-2) C) y - 8 = 4/5 (x-6)

OpenStudy (anonymous):

\[slope = \frac{ y_2 - y_1}{ x_2 - x_1 }\] Point slope form \[y - y_1 = m(x - x_1)\] (x1, y1) = given points m = slope